OpenStudy (anonymous):

lim (1/n^2 + 2/n^2 + ... + n-1/n^2) . n->infinity . n is Z .

OpenStudy (anonymous):



OpenStudy (aravindg):

james help me

OpenStudy (anonymous):

denominators are all \[n^2\]? if so you can add these up.

OpenStudy (anonymous):

get \[\frac{1+2+...+(n-1)}{n^2}\] and the numerator sums to \[\frac{n(n-1)}{2}\] so you should be in good shape to find the limit

OpenStudy (anonymous):

btw, n\in \mathbb Z translates as \[ n \in \mathbb Z\]

OpenStudy (anonymous):

Yes , I know that . Thanks !

OpenStudy (anonymous):

yw